OpenStudy (anonymous):

what is the slope-intercept: equation of the line that passes through the point (6, 4) and is parallel to y = 1/3 x - 3. help me find the answer or the procedure pls

OpenStudy (anonymous):

Well when you have the equation of one line and you want a parrllel lin the slope is going to be the same for both equations. then use the equation y-y1=m(x-x1) to find the anwer. if you post your answer when you get it I'll check it for you.

OpenStudy (anonymous):

so i have to plot the x & y-intercept in the equation but i still dont get how

OpenStudy (anonymous):

let me show you how to find your new equation. so the formula we are going to use to find the ne equation is y-y1=m(x-x1) our slope (m) is the same as the equation givin. so the slope is 1/3 then we are given the point (6,4) this will be what we put in the place of y1 and x1 remember that 6 is x1 and 4 is y1 this time. so y-4=1/3(x-6) now you are going to muliply 1/3 to x and then 1/3 to 6 you should get y-4=1/3x-2 your last step is to add 4 to both sides so that you get 1 by itself. your final answer should be y=1/3x+2

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

oh ok so the equation to find the slope-intercept is y-y1=m(x-x1) right?

OpenStudy (anonymous):

only when you have the slope and at least one set of ordered pairs

OpenStudy (anonymous):

oh ok then thank you very much

OpenStudy (anonymous):

you're welcome